# A general approach to modeling environmental mixtures with multivariate outcomes

### v1

@@ Acknowledgements
Appendix A Model Specification
@@
$y_{iK}$ $\approx\beta_{K0}+\mathbf{b}_{i1,\bm{\theta}_{K1}}^{T}\bm{\beta}_{K1}+\cdots+\mathbf{b}_{iP,\bm{\theta}_{kj}}^{T}\bm{\beta}_{kj}+\bm{z}_{i}^{T}\bm{\beta}_{ZK}+\xi\sigma_{K}u_{i}+\epsilon_{iK}$
$\epsilon_{ik}$ $\overset{iid}{\sim}N(0,\sigma^{2})$
$u_{ik}$ $\overset{iid}{\sim}N(0,1)$
@@
$\mathcal{L}$ $=\prod_{k=1}^{K}\prod_{i=1}^{n}f(y_{ik};\bm{\beta}_{k1},\dots,\bm{\beta}_{kj},\bm{\theta}_{k1},\dots,\bm{\theta}_{kj},u_{i})$
[…]
@@
or $=\prod_{k=1}^{K}\left[\prod_{i=1}^{n}f(y_{ik};\bm{\beta}^{*}_{Z^{\beta}_{k1}},\dots,\bm{\beta}^{*}_{Z^{\beta}_{kj}},\bm{\theta}^{*}_{Z^{\theta}_{k1}},\dots,\bm{\theta}^{*}_{Z^{\theta}_{kj}},u_{i})\right]$
@@
Appendix B MCMC Sampler
@@
    $P(Z^{\beta}_{kj}=a|...)$ $=\frac{\pi_{a,Z^{\theta}_{kj}}\tilde{f}^{\beta}(a)}{\sum_{c=1}^{C}\pi_{c,Z^{\theta}_{kj}}\tilde{f}^{\beta}(c)},~{}~{}a\in\{1,\dots,C\},$
    where $\text{where }\tilde{f}^{\beta}(a)$ $=\left[\prod_{i=1}^{n}f(y_{ik};\bm{\beta}^{*}_{Z^{\beta}_{k1}},\dots,\bm{\beta}^{*}_{Z^{\beta}_{k(j-1)}},\bm{\beta}^{*}_{a},\bm{\beta}^{*}_{Z^{\beta}_{k(j+1)}},\dots,\bm{\beta}^{*}_{Z^{\beta}_{kj}},\bm{\theta}^{*}_{Z^{\theta}_{k1}},\dots,\bm{\theta}^{*}_{Z^{\theta}_{kj}},u_{i})\right]$
    $P(Z^{\theta}_{kj}=b|...)$ $=\frac{\pi_{Z^{\beta}_{kj},b}\tilde{f}^{\theta}(b)}{\sum_{c=1}^{C}\pi_{Z^{\beta}_{kj},b}\tilde{f}^{\theta}(c)},~{}~{}b\in\{1,\dots,C\}$
    where $\text{where }\tilde{f}^{\theta}(b)$ $=\left[\prod_{i=1}^{n}f(y_{ik};\bm{\beta}^{*}_{Z^{\beta}_{k1}},\dots,\bm{\beta}^{*}_{Z^{\beta}_{kj}},\bm{\theta}^{*}_{Z^{\theta}_{k1}},\dots,\bm{\theta}^{*}_{Z^{\theta}_{k(j-1)}},\bm{\theta}^{*}_{b},\bm{\theta}^{*}_{Z^{\theta}_{k(j+1)}},\dots,\bm{\theta}^{*}_{Z^{\theta}_{kj}},u_{i})\right]$
@@
    $(V^{\beta}_{c}|...)$ $\propto(1-V^{\beta}_{c})^{\alpha_{\beta}-1}\prod_{k=1}^{K}\prod_{j=1}^{p}\pi^{*}_{z^{\beta}_{kj},z^{\theta}_{kj}}$
    $(V^{\theta}_{c}|...)$ $\propto(1-V^{\theta}_{c})^{\alpha_{\theta}-1}\prod_{k=1}^{K}\prod_{j=1}^{p}\pi^{*}_{z^{\beta}_{kj},z^{\theta}_{kj}}.$
@@
      $(\phi^{*}_{cj}|...)$ $\propto\exp\left(\tau_{\theta}\mathbf{1}^{T}\bm{\theta}^{*}_{c}-\frac{1}{2}\lambda^{\theta}{\bm{\theta}_{c}^{*}}^{T}\Sigma_{\theta}^{-1}\bm{\theta}_{c}^{*}\right)$
      ×$\qquad\qquad\times\prod_{k=1}^{K}\left[\prod_{i=1}^{n}f(y_{ik};\bm{\beta}^{*}_{Z^{\beta}_{k1}},\dots,\bm{\beta}^{*}_{Z^{\beta}_{kj}},\bm{\theta}^{*}_{Z^{\theta}_{k1}},\dots,\bm{\theta}^{*}_{Z^{\theta}_{kj}},u_{i})\right]^{I(\sum_{j}z^{\theta}_{kj}>0)}$
      ×$\qquad\qquad\times|\cos(\phi^{*}_{cj})^{L-j}|$
@@
      $(\bm{\theta}^{*}_{c}|...)$ $\propto I(||\bm{\theta}^{*}_{c}||=1)\exp\left(\left[\tau_{\theta}\mathbf{1}+\sum_{k,j:z^{\theta}_{kj}=c}\sigma^{-2}_{k}\hat{\mathbf{X}}^{T}_{k}\hat{\mathbf{y}}_{k}\right]^{T}\bm{\theta}^{*}_{c}\right.$
      $\qquad\qquad\qquad\qquad\qquad-\frac{1}{2}{\bm{\theta}^{*}_{c}}^{T}\left[\sum_{k,j:z^{\theta}_{kj}=c}\sigma^{-2}_{k}\hat{\mathbf{X}}_{k}^{T}\hat{\mathbf{X}}_{k}\right]\bm{\theta}^{*}_{c}-\frac{1}{2}\lambda^{\theta}{\bm{\theta}_{c}^{*}}^{T}\Sigma_{\theta}^{-1}\bm{\theta}_{c}^{*}\Biggr{)}$
@@
      In practice sampling from this distribution can be challenging, so we approximate it by sampling from $N\left(\hat{\bm{\mu}},~{}~{}\hat{\bm{\Sigma}}\right)$ and then standardizing appropriately:
@@
    $(\log(\rho)|...)$ $\propto Gamma(a_{\rho},b_{\rho})\times\prod_{k=1}^{K}\prod_{j=1}^{p}\pi^{*}_{z^{\beta}_{kj},z^{\theta}_{kj}}\times\rho$
@@
    $(\text{log}(\xi)|...)$ $\propto InvGamma(a_{\xi},b_{\xi})\times\exp\left(-\frac{1}{2}\sum_{k=1}^{K}\frac{1}{\sigma^{2}_{k}}\sum_{i=1}^{n}\left[y_{ik}-\beta_{0k}-\sum_{j=1}^{p}\mathbf{b}_{i,\bm{\theta}_{kj}}^{T}\bm{\beta}_{kj}-\xi\sigma_{k}u_{i}\right]^{2}\right)\times\xi$
@@
    $(\text{log}(\sigma^{2}_{k})|...)$ $\propto InvGamma\left(a_{\sigma},b_{\sigma}\right)\times(\sigma^{2}_{k})^{-\frac{n}{2}}\exp\left(-\frac{1}{2\sigma^{2}_{k}}\sum_{i=1}^{n}\left[y_{ik}-\beta_{0k}-\sum_{j=1}^{p}\mathbf{b}_{i,\bm{\theta}_{kj}}^{T}\bm{\beta}_{kj}-\xi\sigma_{k}u_{i}\right]^{2}\right)\times\sigma^{2}_{k}$
@@
    $\beta_{0k}|...$ $\sim N\left(\frac{1}{n}\sum_{i=1}^{n}\left[y_{ik}-\sum_{j=1}^{p}\mathbf{b}_{i,\bm{\theta}_{kj}}^{T}\bm{\beta}_{kj}-\xi\sigma_{k}u_{i}\right]^{2},~{}~{}\sigma^{2}_{k}/n\right)$
@@
Appendix C Variable Importance Metrics for Arbitrary Groups of Exposures
When interest lies not in individual exposures but in an important group of exposures $\bm{x}^{\dagger}$—for example a class of highly correlated exposures—our proposed importance measure $\Psi_{kj}$ generalizes to any partition $\bm{x}^{T}=[{\bm{x}^{\dagger}}^{T},{\bm{x}^{o}}^{T}]$:
@@ Appendix C Variable Importance Metrics for Arbitrary Groups of Exposures
Appendix D Additional simulation results
D.1 Illustration of Autocorrelated Exposures for Simulation A
@@
Here we provide additional information on the data generating process for the simulations involving distributed lag exposures in Section 5.1. Figure D7 shows three realizations (three observations) of the exposure vectors measured across time. We see that there is a moderate degree of temporal dependence due to the autoregressive structure with a large correlation coefficient. Additionally, the five exposures tend to increase (decrease) together over time due to the dependence induced by the vector autoregressive process generating the data, which imposes correlation across exposures at each point in time.
@@
Figure D7: Exposure trajectories over time for three randomly chosen observations from one of the simulated data sets in Section 5.1.
D.2 Additional results
@@
Here we provide additional results for the simulation studies found in Section 5.1. Specifically, we now provide 90% credible interval coverage for the $\bm{\omega}_{kj}$ parameters (Figure D9), estimates of the individual $\bm{\omega}_{kj}$ trajectories across data sets and their corresponding means (Figure D10), and mean squared error of estimation of $f_{kj}(\cdot)$ (Figure D8). Overall, the results echo those seen in Section 5.1, but provide some additional insight into clustering, and when it is beneficial. We see that the interval coverages are generally at the desired 90% level, however, for outcome 1 under both exposure 2 and 3, we see some coverage below the nominal level. We see that this issue is reduced when we do not perform clustering as the DLNM clustering estimator has credible interval closer to the nominal level. Figure D10 provides intuition for this, as we can see that in many data sets, exposures 2 and 3 are placed into the wrong cluster for $\bm{\omega}_{kj}$. This generally highlights the bias-variance trade-off that comes from our clustering algorithm. Clustering generally improves overall MSE of these parameters, which are difficult to estimate in small sample sizes, but the shrinkage across exposures and outcomes induces some bias that can lead to reduced credible interval coverages. Lastly, when estimating $f_{kj}(\cdot)$, we see similar results as for $\bm{\omega}_{kj}$, though there are certain instances where it is better to not perform clustering across outcomes, as seen in the top panel of Figure D8. This issue, along with the reduced credible intervals for certain combinations of exposures and outcomes, disappears for larger sample sizes, which we show in the following section.
@@
Figure D8: Mean squared error for estimating $f_{kj}(\cdot)$ across all estimators. MSE is scaled so that the DLNM estimator has an MSE of 1 in all scenarios.
@@
Figure D9: 90% credible interval coverage for estimating $\bm{\omega}_{kj}$ across all estimators.
@@
Figure D10: Estimates of $\bm{\omega}_{kj}$ for the DLNM estimator. Individual grey lines correspond to individual data sets, while the red line represents the truth, and the black line represents the mean across data sets.
D.3 Larger sample sizes

### v2

@@ Acknowledgements
Appendix A Special Cases
Choice of $\mathbf{A}_{j}$ determines the model structure, and we briefly outline here several examples.
  1. 1.
    Suppose for each unit $i$ we observe each of $P$ exposures at $L$ times; that is, $\mathbf{x}_{ip}=[x_{ip1},\dots,x_{ipL}]^{T}$ for $p=1,\dots,P$. We write the full $PL$-vector of exposures $\mathbf{x}_{i}^{*}=\left[\mathbf{x}_{i1}^{T},\dots,\mathbf{x}_{iP}^{T}\right]^{T}$. Setting $J$ to be equal to the number of distinct exposures ($J=P$) and choosing $A_{kj}=\left[\mathbf{0}_{L\times(j-1)L}~|~\mathbf{I}_{L\times L}~|~\mathbf{0}_{L\times(P-j)L}\right]$ (the $L\times PL$ horizontal block matrix with $j^{th}$ block equal to the identity matrix, and all other zero), one obtains a (multivariate) DLNM. We elaborate on this further in the main text, including specification of smoothness inducing centering distribution $G_{0,\theta}$.
  2. 2.
    Suppose for each unit $i$ we observe each of $P$ exposures at $B$ biomarkers. For example, bauer2020associations analyzed five different biomarkers (in hair, blood, urine, nails, and saliva) of four different metals ($P=4$, $B=5$). Let $\mathbf{x}_{ip}=[x_{ip1},\dots,x_{ipB}]^{T}$ for $p=1,\dots,P$. We write the full $PB$-vector of exposures $\mathbf{x}_{i}^{*}=\left[\mathbf{x}_{i1}^{T},\dots,\mathbf{x}_{iP}^{T}\right]^{T}$. Setting $J$ to be equal to the number of distinct exposures ($J=P$) and choosing $A_{kj}=\left[\mathbf{0}_{B\times(j-1)B}~|~\mathbf{I}_{B\times B}~|~\mathbf{0}_{B\times(P-j)B}\right]$ (the $B\times PB$ horizontal block matrix with $j^{th}$ block equal to the identity matrix, and all other zero). This is analogous to the DLNM above, however we would replace the smoothness inducing centering distribution $G_{0,\theta}$ with a default ridge penalty prior.
  3. 3.
    Suppose now we observe $P$ exposures at a single time point, and let $\mathbf{x}_{i}^{*}=\left[{x}_{i1},\dots,{x}_{iP}\right]^{T}$. Setting $J=P$ and $\mathbf{A}_{j}=\mathbf{e}_{j}$ (i.e. the $P$-vector of zeroes except for the $j^{th}$ element, equal to to one) yields a (multivariate) additive model:
    $y_{ik}$ $=f_{k1}({{x}}_{i1})+\cdots+f_{kP}({{x}}_{iP})+\bm{z}_{i}^{T}\bm{\beta}_{Zk}+\xi\sigma_{k}u_{i}+\epsilon_{ik},~~k=1,\dots,K.$
    Here the weights are fixed ($\omega_{kp}=1$), but we can still improve performance by exploiting similarity in the functions $f_{kp}(\cdot)$ across similar exposure-outcome pairs. Even with univariate outcomes ($K=1$) this encourages borrowing strength across multiple exposure-response curves.
  4. 4.
    Suppose again we observe $P$ exposures at a single time point, and let $\mathbf{x}_{i}^{*}=\left[{x}_{i1},\dots,{x}_{iP}\right]^{T}$. Setting $J=1$ and $\mathbf{A}_{1}=\mathbf{I}_{P\times P}$ yields a multivariate single index model (SIM).
    for $k=1,\dots,K$. Cluster-inducing priors allow borrowing information across different outcomes here, but not exposures.
Appendix B Model Specification
@@
$y_{iK}$ $\approx\beta_{K0}+\mathbf{b}_{i1,\bm{\theta}_{K1}}^{T}\bm{\beta}_{K1}+\cdots+\mathbf{b}_{iP,\bm{\theta}_{KP}}^{T}\bm{\beta}_{KP}+\bm{z}_{i}^{T}\bm{\beta}_{ZK}+\xi\sigma_{K}u_{i}+\epsilon_{iK}$
$\epsilon_{ik}$ $\overset{iid}{\sim}N(0,\sigma^{2}_{k})$
$u_{i}$ $\overset{iid}{\sim}N(0,1)$
@@
$\mathcal{L}$ $=\prod_{k=1}^{K}\prod_{i=1}^{n}f(y_{ik};\bm{\beta}_{k1},\dots,\bm{\beta}_{kP},\bm{\theta}_{k1},\dots,\bm{\theta}_{kjP},u_{i})$
$\text{ where }~~\bm{\beta}_{kj}$ $=\sum_{c=1}^{C}I(Z^{\beta}_{kj}=c)\bm{\beta}^{*}_{c}\text{ and }\bm{\theta}_{kj}=\sum_{c=1}^{C}I(Z^{\theta}_{kj}=c)\bm{\theta}^{*}_{c}$
@@
or $=\prod_{k=1}^{K}\left[\prod_{i=1}^{n}f(y_{ik};\bm{\beta}^{*}_{Z^{\beta}_{k1}},\dots,\bm{\beta}^{*}_{Z^{\beta}_{kP}},\bm{\theta}^{*}_{Z^{\theta}_{k1}},\dots,\bm{\theta}^{*}_{Z^{\theta}_{kP}},u_{i})\right]$
@@
Appendix C Posterior sampling of index parameters
Sampling from the full conditional for the index weights $\bm{\theta}^{*}_{c}$ is complicated by its constrained parameter space. We present two approaches to sampling $\bm{\theta}^{*}_{c}$: the first transforms $\bm{\theta}^{*}_{c}$ to an unconstrained space; the second exploits a Taylor approximation to obtain a closed form for the full conditional distribution and ignores the sign constraint.
C.1 Polar Coordinate Reparameterization
The constraints on $\bm{\theta}^{*}_{c}$ makes choosing reasonable proposal distributions challenging in practice. One path forward is to absorb the constraints into the parameterization by transforming to polar coordinates. Following the approach of park2005bayesian, we let
C.2 Approximate Sampling via Taylor Expansion
The approach of Section C.1 can be slow to mix in high-dimensional settings. Alternatively, we propose the following approximate method, ignoring the sign constraint and using the Fisher-Bingham distribution on the full hypersphere for $G_{0,\theta}$. A first order Taylor expansion of $f_{kj}(\tilde{\mathbf{x}}_{ij}^{T}\bm{\theta}^{*}_{c})$ around the previous value $\tilde{\bm{\theta}}^{*}_{c}$ implies that the sum of squares is approximately
where $\hat{\mathbf{y}}_{k}$ is a vector with $i^{th}$ element equal to
This corresponds to a Fisher Bingham distribution, $\bm{\theta}^{*}_{c}\sim FB({\bm{\eta}}^{T},-\frac{1}{2}\bm{\Lambda}^{-1})$, which can be conveniently sampled from via the simdd package in R (kent2018new). Despite its closed form, directly sampling from a Fisher-Bingham distribution can be computationally challenging in high dimensions. Absent the indicator function, the density in (5) is proportional to a multivariate normal distribution for $\bm{\theta}^{*}_{c}$. As such, in high-dimensional settings we propose to use an approximate update, where we first draw a random variable from a $N\left(\bm{\Lambda}{\bm{\eta}},~\bm{\Lambda}\right)$ distribution and then project this value to the unit sphere. Empirically we have found that this approximation leads to very similar samples from the posterior distribution of $\bm{\theta}^{*}_{c}$ and is significantly faster to compute.
Appendix D MCMC Sampler
@@
[…]
    $\text{where }\tilde{f}^{\beta}(a)$ $=\left[\prod_{i=1}^{n}f(y_{ik};\bm{\beta}^{*}_{Z^{\beta}_{k1}},\dots,\bm{\beta}^{*}_{Z^{\beta}_{k(j-1)}},\bm{\beta}^{*}_{a},\bm{\beta}^{*}_{Z^{\beta}_{k(j+1)}},\dots,\bm{\beta}^{*}_{Z^{\beta}_{kP}},\bm{\theta}^{*}_{Z^{\theta}_{k1}},\dots,\bm{\theta}^{*}_{Z^{\theta}_{kP}},u_{i})\right]$
[…]
    $\text{where }\tilde{f}^{\theta}(b)$ $=\left[\prod_{i=1}^{n}f(y_{ik};\bm{\beta}^{*}_{Z^{\beta}_{k1}},\dots,\bm{\beta}^{*}_{Z^{\beta}_{kP}},\bm{\theta}^{*}_{Z^{\theta}_{k1}},\dots,\bm{\theta}^{*}_{Z^{\theta}_{k(j-1)}},\bm{\theta}^{*}_{b},\bm{\theta}^{*}_{Z^{\theta}_{k(j+1)}},\dots,\bm{\theta}^{*}_{Z^{\theta}_{kP}},u_{i})\right]$
@@
[…]
@@
[…]
      $\qquad\qquad\times\prod_{k=1}^{K}\left[\prod_{i=1}^{n}f(y_{ik};\bm{\beta}^{*}_{Z^{\beta}_{k1}},\dots,\bm{\beta}^{*}_{Z^{\beta}_{kP}},\bm{\theta}^{*}_{Z^{\theta}_{k1}},\dots,\bm{\theta}^{*}_{Z^{\theta}_{kP}},u_{i})\right]^{I(\sum_{j}z^{\theta}_{kj}>0)}$
      $\qquad\qquad\times|\cos(\phi^{*}_{cj})^{L-j}|$
@@
[…]
      $\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt-\frac{1}{2}{\bm{\theta}^{*}_{c}}^{T}\left[\sum_{k,j:z^{\theta}_{kj}=c}\sigma^{-2}_{k}\hat{\mathbf{X}}_{k}^{T}\hat{\mathbf{X}}_{k}\right]\bm{\theta}^{*}_{c}-\frac{1}{2}\lambda^{\theta}{\bm{\theta}_{c}^{*}}^{T}\Sigma_{\theta}^{-1}\bm{\theta}_{c}^{*}\Biggr)$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
    $\beta_{0k}|$... $\sim N\left(\frac{1}{n}\sum_{i=1}^{n}\left[y_{ik}-\sum_{j=1}^{p}\mathbf{b}_{i,\bm{\theta}_{kj}}^{T}\bm{\beta}_{kj}-\xi\sigma_{k}u_{i}\right],~~\sigma^{2}_{k}/n\right)$
@@
Appendix E Default Hyperparameter Settings
Users may provide specify their own hyperparameters using the provided software. Below we detail the defaults used in our simulations and analyses unless otherwise stated.
  • Priors hyperparameters:
    • $a_{\beta}=b_{\beta}=1$
    • $a_{\theta}=b_{\theta}=1$
    • $a_{\rho}=b_{\rho}=1$
    • $a_{\lambda}^{\beta}=b_{\lambda}^{\beta}=1$
    • $a_{\lambda}^{\theta}=1$, $b_{\lambda}^{\theta}=0.001$
    • $a_{\xi}=b_{\xi}=0.01$
    • $a_{\sigma}=b_{\sigma}=0.01$
  • MCMC settings:
    • Random walk SD=1 for $\log(\rho)$, $\log(\lambda^{\theta})$, $\xi$, $\log(\sigma^{2}_{k})$
    • Gridsize=10 for grid sampling $V_{c}^{\beta}$ and $V_{c}^{\theta}$
  • Simulation A: Distributed lag models with multiple exposures
    • In models (1) & (2) we set $C=10$ and in model (4) $C=5$. In model (3) there is no clustering.
    • Order of difference matrix ($D_{s}$) $s=2$
    • Dimension of basis $\bm{\Psi}$ is 6 (when using dimension reduction technique in S2.1)
  • Simulation B: Single time point mixture modeling
    • Order of index model $P=3$
    • Maximum number of clusters $C=KP$
  • NMMAPS Data Analysis
    • Maximum number of clusters $C=6$
    • $a_{\lambda}^{\theta}=1$, $b_{\lambda}^{\theta}=0.001$
Appendix F Exposure Importance Metrics: Details and Extension to Exposure Groups
When interest lies not in individual exposures but in a group of exposures $\bm{x}^{\dagger}$—for example a class of highly correlated exposures—our proposed importance measure $\Psi_{kj}$ generalizes to any partition $\bm{x}^{T}=[{\bm{x}^{\dagger}}^{T},{\bm{x}^{o}}^{T}]$:
@@ Appendix F Exposure Importance Metrics: Details and Extension to Exposure Groups
The proposed $\Psi_{kj}$ has connections to a number of related variable importance metrics in the literature. It is closely related to those proposed in shin2024treatment, which instead target effect modification and identify variables that modify associations the most. It is also related to metrics used in nonparametric regression such as Shapley values (shapley1953value) or leave-one-covariate-out (LOCO; lei2018distribution) that target the extent to which mean squared error is reduced when a covariate is omitted. One important distinction is that our proposed metric does not require fitting multiple models (e.g., omitting exposure $j$). Rather, it uses a single fit along with an estimate of the distribution of the exposures. While the distribution of the exposures can be misspecified, we view $\Psi_{kj}$ as a complementary piece of information to aid interpretation, even if the exposure distribution is imperfectly estimated.
Appendix G Additional simulation details
G.1 Details on Autocorrelated Exposures for Simulation A
We drew exposures at the first time point $\mathbf{x}_{\cdot 1}=(x_{11},\dots,x_{P1})^{T}$ from a $\bm{\epsilon}_{1}\sim\mathcal{MVN}(\bm{0},\bm{\Sigma})$ and generated subsequent exposures $\mathbf{x}_{\cdot l}=(x_{1l},\dots,x_{Pl})^{T}$ for $l=2,\dots,L$, according to $\mathbf{x}_{\cdot l}=\bm{M}\mathbf{x}_{\cdot l-1}+\bm{\epsilon}_{l}$, where $\bm{\epsilon}_{l}\sim\mathcal{MVN}(\bm{0},\bm{\Sigma})$. We set $\bm{M}=0.85\bm{I}_{P}$ to induce a high degree of correlation over time within exposures, and we set $\bm{\Sigma}_{pp^{\prime}}=0.6^{|p-p^{\prime}|}$ to induce dependence across exposures, as is typical in mixtures of environmental exposures.
@@ G.1 Details on Autocorrelated Exposures for Simulation A
Here we provide additional information on the data generating process for the simulations involving distributed lag exposures in Section 5.1. Figure G7 shows three realizations (three observations) of the exposure vectors measured across time. We see that there is a moderate degree of temporal dependence due to the autoregressive structure with a large correlation coefficient. Additionally, the five exposures tend to increase (decrease) together over time due to the dependence induced by the vector autoregressive process generating the data, which imposes correlation across exposures at each point in time.
@@
Figure G7: Exposure trajectories over time for three randomly chosen observations from one of the simulated data sets in Section 5.1.
G.2 Additional results
@@
Here we provide additional results for the simulation studies found in Section 5.1. Specifically, we now provide 95% credible interval coverage for the $\bm{\omega}_{kj}$ parameters (Figure G9), estimates of the individual $\bm{\omega}_{kj}$ trajectories across data sets and their corresponding means (Figure G10), and mean squared error of estimation of $f_{kj}(\cdot)$ (Figure G8). Overall, the results echo those seen in Section 5.1, but provide some additional insight into clustering, and when it is beneficial. We see that the interval coverages are generally at the desired 95% level, however, for outcome 2 under both exposure 2 and 3, we see some slight under-coverage below the nominal level. We see that this issue is reduced when we do not perform clustering as the DLNM no clustering estimator has credible interval at the nominal level. Figure G10 provides intuition for this, as we can see that in many data sets, exposures 2 and 3 are placed into the wrong cluster for $\bm{\omega}_{kj}$. This generally highlights the bias-variance trade-off that comes from our clustering algorithm. Clustering generally improves overall MSE of these parameters, which are difficult to estimate in small sample sizes, but the shrinkage across exposures and outcomes induces some bias that can lead to reduced credible interval coverages. Lastly, when estimating $f_{kj}(\cdot)$, we see similar results as for $\bm{\omega}_{kj}$, though there are certain instances where it is slightly better to not perform clustering across outcomes, as seen in the bottom panel of Figure G8. This issue, along with the reduced credible interval coverages for certain combinations of exposures and outcomes, disappears for larger sample sizes, which we show in the following section.
@@
Figure G8: Mean squared error for estimating $f_{kj}(\cdot)$ across all estimators. MSE is scaled so that the DLNM estimator has an MSE of 1 in all scenarios.
@@
Figure G9: 95% credible interval coverage for estimating $\bm{\omega}_{kj}$ across all estimators.
@@
Figure G10: Estimates of $\bm{\omega}_{kj}$ for the DLNM estimator. Individual grey lines correspond to individual data sets, while the red line represents the truth, and the black line represents the mean across data sets.
G.3 Larger sample sizes